OpenStudy (anonymous):

please help me: if log a(x+61)=3 and a^3= 125 determine 3^ √x+61

OpenStudy (raden):

solve for a, first a^3= 125 a^3= 5^3 a = 5

OpenStudy (raden):

now see the first equation, log a (x+61)=3 put a = 5, get log5 (x+61) = 3 change into exponent form : as x+61 = 5^3 x+61 = 125 x = 125 - 61 x = 64

OpenStudy (raden):

the last, to get the value of 3^ √x+61 subtitute x = 64, get 3^ √64+61 = 3^ √125 = 5

OpenStudy (raden):

hope that make sense :)

OpenStudy (anonymous):

hey thanks

OpenStudy (raden):

welcome